OpenStudy (goformit100):

A and B are two candidates seeking admission in a college. The probability that A is selected is 0.7 and the probability that exactly one of them is selected is 0.6. Find the probability that B is selected.

OpenStudy (goformit100):

@primeralph

OpenStudy (anonymous):

probability that just one is selected=(probability that each one is selected while the other isn't) = (1-A)B+A(1-B)=.6 A is known, so this becomes .3B+.7(1-B)=.6 Solve for B.